.com/hm.tpx 8. 1000 value 10.00 points Problem 18-21 Cash dividend versus stock repurchase [LO18-5) The Carlton Corporation has $6 million in earnings after tax cash es and 3 milion shares outstanding. The stock trades at a P/E of 10. The firm has $2 million in excess a. Compute the current price of the stock. ( Do not round intermediate calculations and round your answer to 2 decimal places.) price b. f the $2 million is used to pay dividends, how much will dividends per share be? (Do not round intermediate calculations and round your answer to 2 decimal places.) Dividends per share
Answer #1

a. Earnings per share = Stock price / Number of shares = $6,000,000 / 3,000,000 = $2.00

Stock price = P/E ratio × Earnings per share = 10 × $2.00 = $20.00

b. Dividends per share = Total dividends / Number of shares = $2,000,000 / 3,000,000 = $0.67